# Simultaneous spatial-parametric collocation approximation for parametric PDEs with log-normal random inputs

### v4

@@ 2.2 Extended least squares sampling algorithms in Bochner spaces
$\varrho_{n}(F,L_{2}(U,{\mathbb{C}};\mu))\ \leq\ \sup_{f\in F}\left\|{f-\tilde{S}_{n}^{\mathbb{C}}f}\right\|_{L_{2}(U,{\mathbb{C}};\mu)}\ \leq\left(\frac{c_{p}}{k}\sum_{s\geq k/8}d_{s}^{p}\right)^{1/p},$
@@
$\tilde{S}_{n}^{\mathbb{C}}:=S_{n}^{\mathbb{C}}({\boldsymbol{y}}_{1},\dots,{\boldsymbol{y}}_{n},\omega_{1},\dots,\omega_{n},\Phi_{k})\quad\text{and}\quad n\geq k\geq\frac{n}{43200}\,.$ (2.21)
@@
$\varrho_{n}(B_{X,{\boldsymbol{\sigma}}},L_{2}(U,X;\mu))\ \leq\ \sup_{v\in B_{X,{\boldsymbol{\sigma}}}}\left\|{v-\tilde{S}_{n}^{X}v}\right\|_{L_{2}(U,X;\mu)}\ \leq\left(\frac{c_{p}}{k}\sum_{s\geq k/8}d_{s}^{p}\right)^{1/p},$
@@
$\tilde{S}_{n}^{X}:=S_{n}^{X}({\boldsymbol{y}}_{1},\dots,{\boldsymbol{y}}_{n},\omega_{1},\dots,\omega_{n},\Phi_{k})\quad\text{and}\quad n\geq k\geq\frac{n}{43200}\,.$ (2.22)
@@
Let $0<p<2$, $m\in{\mathbb{N}}_{0}$ and $d_{s}:=d_{s}(\delta_{m}B_{\mathbb{C}}^{\alpha,\tau},L_{2}(U,{\mathbb{C}};\mu))$. Assume that there is a constant $K$ such that for every $v\in H_{X}^{\alpha,\tau}$ and every $m\in{\mathbb{N}}_{0}$, $\left\|{P_{m}v}\right\|_{H_{X}^{\alpha,\tau}}\leq K\left\|{f}\right\|_{H_{X}^{\alpha,\tau}}$. Then there is a constant $c_{p}\in{\mathbb{N}}$ such that for any $n\in{\mathbb{N}}$, there exist points ${\boldsymbol{y}}_{1},\dots,{\boldsymbol{y}}_{n}\in U_{0}$ and weights $\omega_{1},\dots,\omega_{n}$ such that
$\varrho_{n}(\delta_{m}B_{X}^{\alpha,\tau},L_{2}(U,X;\mu))\ \leq\ \sup_{\delta_{m}v\in\delta_{m}B_{X}^{\alpha,\tau}}\left\|{\delta_{m}v-\tilde{S}_{n}^{X}(\delta_{m}v)}\right\|_{L_{2}(U,X;\mu)}\ \leq\left(\frac{c_{p}}{k}\sum_{s\geq k/8}d_{s}^{p}\right)^{1/p},$
where $\tilde{S}_{n}^{X}$ is defined as in (2.22).
Let $\Lambda\subset{\mathbb{N}}$ be a finite set. For $v\in L_{2}(U,X;\mu)$, we define the truncation of the GPC expansion of $v$
$S_{\Lambda}v:=\sum_{s\in\Lambda}{\color[rgb]{0,0,0}{v_{s}}}\varphi_{s}.$
@@
Assume that there is a constant $K$ such that for every $v\in H_{X}^{\alpha,\tau}$ and every $m\in{\mathbb{N}}_{0}$, $\left\|{P_{m}v}\right\|_{H_{X}^{\alpha,\tau}}\leq K\left\|{v}\right\|_{H_{X}^{\alpha,\tau}}$. For $\xi>0$ and $m\in{\mathbb{N}}_{0}$, we introduce the set
@@
$\|\delta_{m}v-S_{\Lambda(\xi)}(\delta_{m}v)\|_{L_{2}(U,X;\mu)}\ \leq\ 2K\xi^{-2/q_{1}}\|\delta_{m}v\|_{H_{X}^{\alpha,\tau}},\ \ \ v\in H_{X}^{\alpha,\tau}.$ (2.26)
Moreover, for all $n\geq 2$ there exists a number $\xi_{n}$ such that $|\Lambda_{m}(\xi_{n})|\leq n$ , and
$\color[rgb]{0,0,0}{\sup_{\delta_{m}v\in\delta_{m}B_{X}^{\alpha,\tau}}\|\delta_{m}v-S_{\Lambda_{m}(\xi_{n})}(\delta_{m}v)\|_{L_{2}(U,X;\mu)}\ \leq\ C\,(2^{m}n)^{-\beta}(m+\log n)^{\beta\tau/\alpha}\ \ \forall n\in{\mathbb{N}}.}$} (2.27)
Proof. From the assumptions we can see that for every $v\in H_{X}^{\alpha,\tau}$ and every $m\in{\mathbb{N}}_{0}$, $\left\|{\delta_{m}v}\right\|_{H_{X}^{\alpha,\tau}}\leq 2K\left\|{v}\right\|_{H_{X}^{\alpha,\tau}}$. For a function $v\in B_{\mathbb{C}}^{\alpha,\tau}$, putting
@@
$\begin{split}\|(\delta_{m}v)-(\delta_{m}v)_{\xi}\|_{L_{2}(U,X;\mu)}^{2}\ &=\ \sum_{{\color[rgb]{0,0,0}{s\in{\mathbb{N}}:}}\,\sigma_{1;s}>\xi^{1/q_{1}}}\|(\delta_{m}v)_{s}\|_{X}^{2}\ =\ \sum_{{\color[rgb]{0,0,0}{s\in{\mathbb{N}}:}}\,\sigma_{1;s}>\xi^{1/q_{1}}}(\sigma_{1;s}\|(\delta_{m}v)_{s}\|_{X})^{2}\sigma_{1;s}^{-2}\\[6.45831pt] \ &\leq\ \xi^{-2/q_{1}}\sum_{{\color[rgb]{0,0,0}{s\in{\mathbb{N}}:}}\,\sigma_{1;s}>\xi^{1/q_{1}}}(\sigma_{1;s}\|(\delta_{m}v)_{s}\|_{X})^{2}\\[6.45831pt] \ &\leq\ (2K)^{2}\xi^{-2/q_{1}}\|\delta_{m}v\|_{H_{X}^{\alpha,\tau}}^{2}.\end{split}$
@@
$\begin{split}\|(\delta_{m}v)_{\xi}-S_{\Lambda_{m}(\xi)}(\delta_{m}v)\|_{L_{2}(U,X;\mu)}^{2}\ &=\ \sum_{{\color[rgb]{0,0,0}{s\in{\mathbb{N}}:}}\,\sigma_{1;s}\leq\xi^{1/q_{1}},\ \sigma_{2;s}>\xi^{1/q_{1}}2^{-\alpha m}m^{\tau}}\left\|{(\delta_{m}v)_{s}}\right\|_{X}^{2}\\ &\ \leq\ \sum_{{\color[rgb]{0,0,0}{s\in{\mathbb{N}}:}}\,\sigma_{1;s}\leq\xi^{1/q_{1}},\ \sigma_{2;s}>\xi^{1/q_{1}}2^{-\alpha m}m^{\tau}}\left(2^{\alpha m}m^{-\tau}\xi^{-1/q_{1}}\sigma_{2;s}\left\|{(\delta_{m}v)_{s}}\right\|_{X}\right)^{2}\\[6.45831pt] \ &\leq\ \xi^{-2/q_{1}}\sum_{s\in{\mathbb{N}}}\left(2^{\alpha m}m^{-\tau}\sigma_{2;s}\left\|{(\delta_{m}v)_{s}}\right\|_{X}\right)^{2}\\[6.45831pt] \ &\leq\ (2K)^{2}\xi^{-2/q_{1}}\|\delta_{m}v\|_{H_{X}^{\alpha,\tau}}^{2}.\end{split}$
@@
$\begin{split}{\color[rgb]{0,0,0}{\ |\Lambda_{m}(\xi)|}}\ &=\ \sum_{s\in\Lambda_{m}(\xi)}1\ =\ 2^{-m}\sum_{{\color[rgb]{0,0,0}{s\in{\mathbb{N}}:}}\,\sigma_{1;s}\leq\xi^{1/q_{1}},\ \sigma_{2;s}>\xi^{1/q_{1}}2^{-\alpha m}m^{\tau}}2^{m}\\ \ &\leq\ 2^{-m}\sum_{s\in{\mathbb{N}}:\ \sigma_{1;s}\leq\xi^{1/q_{1}}}\ \sum_{{\color[rgb]{0,0,0}{k\in{\mathbb{N}}_{0}:}}\ \sigma_{2;s}>\xi^{1/q_{1}}2^{-\alpha k}k^{\tau}}2^{k}\\ \ &\leq\ 2^{-m}\sum_{\sigma_{1;{\boldsymbol{s}}}\leq\xi^{1/q_{1}}}\xi^{1/(q_{1}\alpha)}\left(\log\xi\right)^{\tau/\alpha}\sigma_{2;{\boldsymbol{s}}}^{-1/\alpha}\\ \ &\leq\ 2^{-m}\xi^{1/(q_{1}\alpha)}\left(\log\xi\right)^{\tau/\alpha}\left(\sum_{{\color[rgb]{0,0,0}{s\in{\mathbb{N}}:}}\,\sigma_{1;{\boldsymbol{s}}}\leq\xi^{1/q_{1}}}\sigma_{2;{\boldsymbol{s}}}^{-q_{2}}\right)^{1/q}\left(\sum_{{\color[rgb]{0,0,0}{s\in{\mathbb{N}}:}}\,\sigma_{1;{\boldsymbol{s}}}\leq\xi^{1/q_{1}}}1\right)^{1/q^{\prime}}\\ \ &\leq 2^{-m}\xi^{1/(q_{1}\alpha)}\left(\log\xi\right)^{\tau/\alpha}\left(\sum_{{\boldsymbol{s}}\in{\mathbb{F}}}\sigma_{2;{\boldsymbol{s}}}^{-q_{2}}\right)^{1/q}\left(\sum_{{\boldsymbol{s}}\in{\mathbb{F}}}\xi\sigma_{1;{\boldsymbol{s}}}^{-q_{1}}\right)^{1/q^{\prime}}\\ \ &\leq 2^{-m}\xi^{1+\delta/\alpha}\left(\log\xi\right)^{\tau/\alpha}.\end{split}$
For any $n\in{\mathbb{N}}$, letting $\xi_{n}$ be a number satisfying the inequalities
@@
we derive that $\dim{\mathcal{V}}(\Lambda_{m}(\xi_{n}))\leq n$. On the other hand, by (2.28),
$\begin{split}\xi_{n}^{-1/q_{1}}\ \leq\ C\,\left(2^{m}n(\log n+m)^{-\tau/\alpha}\right)^{-\frac{1}{q_{1}}\frac{\alpha}{\alpha+\delta}}\ \leq\ C\,(2^{m}n)^{-\beta}(m+\log n)^{\beta\tau/\alpha}.\end{split}$
This together with (2.26) proves (2.27).
@@
$\|\delta_{m}f-S_{\Lambda(\xi)}(\delta_{m}f)\|_{L_{2}(U,{\mathbb{C}};\mu)}\ \leq\ 2K\xi^{-2/q_{1}}\|\delta_{m}f\|_{H_{\mathbb{C}}^{\alpha,\tau}},\ \ \forall f\in H_{\mathbb{C}}^{\alpha,\tau}.$
Moreover, for all $n\geq 2$ there exists a number $\xi_{n}$ such that $\dim({\mathcal{V}}(\Lambda_{m}(\xi_{n}))\leq n$, and
$\begin{split}d_{n}(\delta_{m}B_{\mathbb{C}}^{\alpha,\tau},L_{2}(U,{\mathbb{C}};\mu))\ &\leq\ \sup_{\delta_{m}f\in\delta_{m}B_{\mathbb{C}}^{\alpha,\tau}}\|\delta_{m}f-S_{\Lambda_{m}(\xi_{n})}(\delta_{m}f)\|_{L_{2}(U,{\mathbb{C}};\mu)}\\ \ &\leq\ C\,(2^{m}n)^{-\beta}(m+\log n)^{\beta\tau/\alpha}\ \ \forall n\in{\mathbb{N}}.\end{split}$
@@ 2.3 Multi-level least squares sampling algorithms
Let $0<q_{1}\leq q_{2}<2$ and $p$ be a fixed number such that $1/q_{2}<p<2$. Let the numbers $\beta$ and $\delta$ be defined as in (2.25). Let $\tau$ be any number such that $\tau>1$ if $\alpha\leq 1/q_{2}$, and $\tau=\alpha/(\alpha-\beta)$ if $\alpha>1/q_{2}$. Let $(\xi_{n})_{n\in{\mathbb{N}}}$ be a sequence of increasing positive numbers $\xi_{n}$ satisfying the condition
@@
Assume that there is a constant $K:=K_{\tau}$ such that for every $v\in H_{X}^{\alpha,\tau}$ and every $m\in{\mathbb{N}}_{0}$, $\left\|{P_{m}v}\right\|_{H_{X}^{\alpha,\tau}}\leq K\left\|{v}\right\|_{H_{X}^{\alpha,\tau}}$. Then for all $n\geq 2$ there exist ${\boldsymbol{y}}_{k,1},...,{\boldsymbol{y}}_{k,n_{k}}\in{\color[rgb]{0,0,0}{U_{0}}}$ and $h_{k,1},...,h_{k,n_{k}}\in L_{2}(U,{\mathbb{C}};\mu)$, $k=0,...,k_{n}$, such that for the operator ${\mathcal{S}}_{n}^{X}$ defined as in (2.29)–(2.31),
@@
$\begin{split}\sup_{v\in B_{X}^{\alpha,\tau}}\|v-{\mathcal{S}}^{X}_{n}v\|_{L_{2}(U,X;\mu)}\ \leq\ C_{\tau}\begin{cases}n^{-\alpha}(\log n)^{\tau}\quad&{\rm if}\ \alpha<1/q_{2},\\ n^{-\alpha}(\log n)^{1+\tau}\quad&{\rm if}\ \alpha=1/q_{2},\\ n^{-\beta}(\log n)^{\tau}\quad&{\rm if}\ \alpha>1/q_{2}.\end{cases}\end{split}$
@@
$\begin{split}\|\delta_{k}v-S_{n_{k}}^{X}(\delta_{k}v)\|_{L_{2}(U,X;\mu)}\ &\leq\ C\,2^{-\alpha k}k^{\tau}\left(\frac{1}{n_{k}}\sum_{j\geq n_{k}}j^{-p/q_{2}}\right)^{1/p}\ \leq\ C\,2^{\left(1/q_{2}-\alpha\right)k}k^{\tau}n^{-1/q_{2}}.\end{split}$ (2.36)
@@
By the equality $k_{n}:=\lfloor\log\xi_{n}\rfloor$ and the equivalence $\xi_{n}\asymp n$, from (2.34) and (2.36) we deduce that
@@
$\sum_{k=0}^{k_{n}}\|\delta_{k}v-{\color[rgb]{0,0,0}{S_{n_{k}}^{X}}}(\delta_{k}v)\|_{L_{2}(U,X;\mu)}\ \leq\ C\,n^{-1/q_{2}}\sum_{k=0}^{k_{n}}2^{\left(1/q_{2}-\alpha\right)k}k^{\tau}\ \leq\ C\,n^{-\alpha}(\log n)^{\tau},$
@@
$\sum_{k=0}^{k_{n}}\|\delta_{k}v-{\color[rgb]{0,0,0}{S_{n_{k}}^{X}}}(\delta_{k}v)\|_{L_{2}(U,X;\mu)}\ \leq\ C\,n^{-1/q_{2}}\sum_{k=0}^{k_{n}}k^{\tau}\ \leq\ C\,n^{-\alpha}(\log n)^{1+\tau}.$
@@
$\begin{split}\|v-{\mathcal{S}}_{n}^{X}v\|_{L_{2}(U,X;\mu)}\ &\leq\ C\,n^{-\alpha}(\log n)^{\tau},\ \ \text{if}\ \ \alpha<1/q_{2},\end{split}$
@@
$\begin{split}\|v-{\mathcal{S}}_{n}^{X}v\|_{L_{2}(U,X;\mu)}\ &\leq\ C\,n^{-\alpha}(\log n)^{1+\tau},\ \ \text{if}\ \ \alpha=1/q_{2}.\end{split}$
@@
We now consider the case $\alpha>1/q_{2}$. In this case we have $\alpha>\beta$. For any $n\in{\mathbb{N}}$, choose $\xi_{n}$ as a number satisfying the inequalities
@@
By the equivalence $\xi_{n}\asymp n^{\frac{1}{q_{1}}\frac{1}{\alpha+\delta}}$, from (2.34) we deduce that
@@
$\begin{split}\|\delta_{k}v-S_{n_{k}}^{X}(\delta_{k}v)\|_{L_{2}(U,X;\mu)}\ &\leq\ \left\|{\delta_{k}v}\right\|_{H_{X}^{\alpha,\tau}}\left(\frac{1}{n_{k}}\sum_{j\geq n_{k}}d_{j}^{p}\right)^{1/p},\end{split}$
@@
$\sum_{k=0}^{k_{n}}\|\delta_{k}v-{\color[rgb]{0,0,0}{S_{n_{k}}^{X}}}(\delta_{k}v)\|_{L_{2}(U,X;\mu)}\ \leq\ \ C\,n^{-\beta}(\log n)^{\tau},$
@@
$\begin{split}\|v-{\mathcal{S}}_{n}^{X}v\|_{L_{2}(U,X;\mu)}\ &\leq\ C\,n^{-\beta}(\log n)^{\tau},\ \ \text{if}\ \ \alpha>1/q_{2}.\end{split}$
@@
By the continuous embedding $H_{X}^{\alpha}\hookrightarrow H_{X}^{\alpha,1+\varepsilon}$ for any $\varepsilon>0$, Theorem 2.1 can be reformulated more conveniently for the space $H_{X}^{\alpha}$ as follows.
@@
Let $0<q_{1}\leq q_{2}<2$ and $\varepsilon$ be any positive number. Let the number $\beta$ be defined as in (2.25). Let $(\xi_{n})_{n\in{\mathbb{N}}}$ be a sequence of increasing positive numbers $\xi_{n}$ satisfying the condition (2.32). Assume that there is a constant $K_{\varepsilon}$ such that for every $v\in H_{X}^{\alpha,1+\varepsilon}$ and every $m\in{\mathbb{N}}_{0}$, $\left\|{P_{m}v}\right\|_{H_{X}^{\alpha,1+\varepsilon}}\leq K_{\varepsilon}\left\|{f}\right\|_{H_{X}^{\alpha,1+\varepsilon}}$. Then for all $n\geq 2$ there exist ${\boldsymbol{y}}_{k,1},...,{\boldsymbol{y}}_{k,n_{k}}\in{\color[rgb]{0,0,0}{U_{0}}}$ and $h_{k,1},...,h_{k,n_{k}}\in L_{2}(U,{\mathbb{C}};\mu)$, $k=0,...,k_{\lceil n/\log n\rceil}$, such that for the operator $\bar{{\mathcal{S}}}^{X}_{n}:={\mathcal{S}}_{\lceil n/\log n\rceil}^{X}$ defined as in (2.29)–(2.31),

### v5

@@ 2.2 Extended least squares sampling algorithms in Bochner spaces
[…]
@@
$S_{n}^{\mathbb{C}}:=S_{n}^{\mathbb{C}}({\boldsymbol{y}}_{1},\dots,{\boldsymbol{y}}_{n},\omega_{1},\dots,\omega_{n},\Phi_{k})\quad\text{and}\quad n\geq k\geq\frac{n}{43200}\,.$ (2.20)
@@
$\varrho_{n}(B_{X,{\boldsymbol{\sigma}}},L_{2}(U,X;\mu))\ \leq\ \sup_{v\in B_{X,{\boldsymbol{\sigma}}}}\left\|{v-S_{n}^{X}v}\right\|_{L_{2}(U,X;\mu)}\ \leq\left(\frac{c_{p}}{k}\sum_{s\geq k/8}d_{s}^{p}\right)^{1/p},$
@@
$S_{n}^{X}:=S_{n}^{X}({\boldsymbol{y}}_{1},\dots,{\boldsymbol{y}}_{n},\omega_{1},\dots,\omega_{n},\Phi_{k})\quad\text{and}\quad n\geq k\geq\frac{n}{43200}\,.$ (2.22)
@@
[…]
$\varrho_{n}(\delta_{m}B_{X}^{\alpha,\tau},L_{2}(U,X;\mu))\ \leq\ \sup_{\delta_{m}v\in\delta_{m}B_{X}^{\alpha,\tau}}\left\|{\delta_{m}v-S_{n}^{X}(\delta_{m}v)}\right\|_{L_{2}(U,X;\mu)}\ \leq\left(\frac{c_{p}}{k}\sum_{s\geq k/8}d_{s}^{p}\right)^{1/p},$
where $S_{n}^{X}$ and $k$ are defined as in (2.22).
Let $\Lambda\subset{\mathbb{N}}$ be a finite set. Let
${\mathcal{V}}(\Lambda)=\operatorname{span}\{\varphi_{s}:\,s\in\Lambda\}$
be the subspace of $L_{2}(U,X;\mu)$, spanned by the functions $\varphi_{s}$, $s\in\Lambda$. For $v\in L_{2}(U,X;\mu)$, we define the truncation $S_{\Lambda}v$ belonging to ${\mathcal{V}}(\Lambda)$, of the GPC expansion of $v$ by
$S_{\Lambda}v:=\sum_{s\in\Lambda}{\color[rgb]{0,0,0}{v_{s}}}\varphi_{s}\in{\mathcal{V}}(\Lambda).$
@@
Assume that there is a constant $K$ such that for every $v\in H_{X}^{\alpha,\tau}$ and every $m\in{\mathbb{N}}_{0}$, $\left\|{P_{m}v}\right\|_{H_{X}^{\alpha,\tau}}\leq K\left\|{v}\right\|_{H_{X}^{\alpha,\tau}}$. For $\xi>1$ and $m\in{\mathbb{N}}_{0}$, we introduce the set
@@
[…]
Moreover, for all $n\geq 2$ and $m\in{\mathbb{N}}_{0}$ there exists a number $\xi_{n}$ such that $|\Lambda_{m}(\xi_{n})|\leq n$ , and
$\color[rgb]{0,0,0}{\sup_{v\in B_{X}^{\alpha,\tau}}\|\delta_{m}v-S_{\Lambda_{m}(\xi_{n})}(\delta_{m}v)\|_{L_{2}(U,X;\mu)}\ \leq\ C\,(2^{m}n)^{-\beta}(m+\log n)^{\beta\tau/\alpha},}$} (2.27)
where the positive constant $C$ depends on $\alpha,q_{1},q_{2}$ and $K$ only.
Proof. From the assumptions we can see that for every $v\in H_{X}^{\alpha,\tau}$ and every $m\in{\mathbb{N}}_{0}$,
$\left\|{\delta_{m}v}\right\|_{H_{X}^{\alpha,\tau}}\leq 2K\left\|{v}\right\|_{H_{X}^{\alpha,\tau}}.$
For a function $v\in B_{X}^{\alpha,\tau}$ , putting
@@
[…]
@@
[…]
@@
[…]
For any $n\geq 2$, letting $\xi_{n}$ be a number satisfying the inequalities
@@
we derive that $\dim{\mathcal{V}}(\Lambda_{m}(\xi_{n}))\leq n$, and moreover,
$\log\xi\ \leq\ m+\log n.$
From the last inequality and the second inequality in (2.28) it follows that
$\begin{split}\xi_{n}^{-1/q_{1}}\ \leq\ C\,\left(2^{m}n(\log n+m)^{-\tau/\alpha}\right)^{-\frac{1}{q_{1}}\frac{\alpha}{\alpha+\delta}}\ \leq\ C\,(2^{m}n)^{-\beta}(m+\log n)^{\beta\tau/\alpha},\end{split}$
where $C$ is a positive constant depending on on $\alpha,q_{1},q_{2}$ only. This together with (2.26) proves (2.27).
@@
[…]
Moreover, for all $n\geq 2$ and $m\in{\mathbb{N}}_{0}$ there exists a number $\xi_{n}$ such that $\dim({\mathcal{V}}(\Lambda_{m}(\xi_{n}))\leq n$, and
$\begin{split}d_{n}(\delta_{m}B_{\mathbb{C}}^{\alpha,\tau},L_{2}(U,{\mathbb{C}};\mu))\ &\leq\ \sup_{{\color[rgb]{0,0,0}{f\in B_{\mathbb{C}}^{\alpha,\tau}}}}\|\delta_{m}f-S_{\Lambda_{m}(\xi_{n})}(\delta_{m}f)\|_{L_{2}(U,{\mathbb{C}};\mu)}\\ \ &\leq\ C\,(2^{m}n)^{-\beta}(m+\log n)^{\beta\tau/\alpha},\end{split}$
where the positive constant $C$ depends on $\alpha,\tau,q_{1},q_{2}$ and $K$ only.
@@ 2.3 Multi-level least squares sampling algorithms
Let $0<q_{1}\leq q_{2}<2$ and $p$ be a fixed number such that $1/q_{2}<p<2$. Let the numbers $\beta$ and $\delta$ be defined as in (2.25). Let $\tau$ be any fixed number such that $\tau>1$ if $\alpha\leq 1/q_{2}$, and $\tau=\alpha/(\alpha-\beta)$ if $\alpha>1/q_{2}$. Let $(\xi_{n})_{n\in{\mathbb{N}}}$ be a sequence of increasing positive numbers $\xi_{n}$ satisfying the condition
@@
Assume that there is a constant $K$ such that for every $v\in H_{X}^{\alpha,\tau}$ and every $m\in{\mathbb{N}}_{0}$,
$\left\|{P_{m}v}\right\|_{H_{X}^{\alpha,\tau}}\leq K\left\|{v}\right\|_{H_{X}^{\alpha,\tau}}.$
Then for all $n\geq 2$ there exist ${\boldsymbol{y}}_{k,1},...,{\boldsymbol{y}}_{k,n_{k}}\in{\color[rgb]{0,0,0}{U_{0}}}$ and $h_{k,1},...,h_{k,n_{k}}\in L_{2}(U,{\mathbb{C}};\mu)$, $k=0,...,k_{n}$, such that for the operator ${\mathcal{S}}_{n}^{X}$ defined as in (2.29)–(2.31),
@@
$\begin{split}\sup_{v\in B_{X}^{\alpha,\tau}}\|v-{\mathcal{S}}^{X}_{n}v\|_{L_{2}(U,X;\mu)}\ \leq\ \begin{cases}C_{\tau}n^{-\alpha}(\log n)^{\tau}\quad&{\rm if}\ \alpha<1/q_{2},\\ C_{\tau}n^{-\alpha}(\log n)^{1+\tau}\quad&{\rm if}\ \alpha=1/q_{2},\\ Cn^{-\beta}(\log n)^{\alpha/(\alpha-\beta)}\quad&{\rm if}\ \alpha>1/q_{2},\end{cases}\end{split}$
where the positive constant $C$ depends only on $\alpha,q_{1},q_{2}$ and $K$; the positive constants $C_{\tau}$ depend only on $\alpha,q_{1},q_{2},K,$ and, explicitly, on $\tau$.
@@
$\begin{split}\|\delta_{k}v-S_{n_{k}}^{X}(\delta_{k}v)\|_{L_{2}(U,X;\mu)}\ &\leq\ C\,2^{-\alpha k}k^{\tau}\left(\frac{1}{n_{k}}\sum_{j\geq n_{k}}j^{-p/q_{2}}\right)^{1/p}\ \leq\ C^{*}\,2^{\left(1/q_{2}-\alpha\right)k}k^{\tau}n^{-1/q_{2}},\end{split}$ (2.36)
where $C^{*}$ is a positive constant depending on $q_{2}$ and $p$ only.
@@
With this choice from the equality $k_{n}:=\lfloor\log\xi_{n}\rfloor$, (2.34) and (2.36) we deduce that
@@
$\sum_{k=0}^{k_{n}}\|\delta_{k}v-{\color[rgb]{0,0,0}{S_{n_{k}}^{X}}}(\delta_{k}v)\|_{L_{2}(U,X;\mu)}\ \leq\ C^{*}\,n^{-1/q_{2}}\sum_{k=0}^{k_{n}}2^{\left(1/q_{2}-\alpha\right)k}k^{\tau}\ \leq\ {\color[rgb]{0,0,0}{C^{\prime}_{\tau}\,n^{-\alpha},}}$
with
$C^{\prime}_{\tau}:=C^{*}\sum_{k=0}^{\infty}2^{-\left(\alpha-1/q_{2}\right)k}k^{\tau},$
@@
$\begin{split}\sum_{k=0}^{k_{n}}\|\delta_{k}v-{\color[rgb]{0,0,0}{S_{n_{k}}^{X}}}(\delta_{k}v)\|_{L_{2}(U,X;\mu)}\ &\leq\ C^{*}\,n^{-1/q_{2}}\sum_{k=0}^{k_{n}}k^{\tau}\\ \ &{\color[rgb]{0,0,0}{\leq\ C^{*}\,n^{-\alpha}\int_{0}^{k_{n}}x^{\tau}{\rm d}x\ \leq\ C^{{}^{\prime\prime}}_{\tau}\,n^{-\alpha}(\log n)^{1+\tau},}}\end{split}$
with
$C^{{}^{\prime\prime}}_{\tau}:=C^{*}(\tau+1)^{-1}.$
@@
$\begin{split}\|v-{\mathcal{S}}_{n}^{X}v\|_{L_{2}(U,X;\mu)}\ &\leq\ {\color[rgb]{0,0,0}{C_{\tau}}}\,n^{-\alpha}(\log n)^{\tau}\ \ {\color[rgb]{0,0,0}{\text{with}\ \ C_{\tau}:=\max\left(2^{\alpha},C^{\prime}_{\tau}\right)}}\ \ \text{if}\ \ \alpha<1/q_{2},\end{split}$ (2.37)
@@
$\begin{split}\|v-{\mathcal{S}}_{n}^{X}v\|_{L_{2}(U,X;\mu)}\ &\leq\ {\color[rgb]{0,0,0}{C_{\tau}}}\,n^{-\alpha}(\log n)^{1+\tau}\ \ {\color[rgb]{0,0,0}{\text{with}\ \ C_{\tau}:=\max\left(2^{\alpha},C^{{}^{\prime\prime}}_{\tau}\right)}}\ \ \text{if}\ \ \alpha=1/q_{2}.\end{split}$ (2.38)
@@
We now consider the case $\alpha>1/q_{2}$. In this case, we have $\alpha>\beta$. For any $n\in{\mathbb{N}}$, choose $\xi_{n}$ as a number satisfying the inequalities
@@
From the last inequalities and (2.34) we deduce that
@@
[…]
@@
$\sum_{k=0}^{k_{n}}\|\delta_{k}v-{\color[rgb]{0,0,0}{S_{n_{k}}^{X}}}(\delta_{k}v)\|_{L_{2}(U,X;\mu)}\ \leq\ \ C\,n^{-\beta}(\log n)^{\tau}.$
@@
$\begin{split}\|v-{\mathcal{S}}_{n}^{X}v\|_{L_{2}(U,X;\mu)}\ &\leq\ C\,n^{-\beta}{\color[rgb]{0,0,0}{(\log n)^{\alpha/(\alpha-\beta)}}},\ \ \text{if}\ \ \alpha>1/q_{2}.\end{split}$
@@
By the continuous embedding $H_{X}^{\alpha}\hookrightarrow H_{X}^{\alpha,1+\varepsilon}$ for any $\varepsilon>0$, taking $\varepsilon=1$ for the case $\alpha<1/q_{2}$, Theorem 2.1 can be reformulated more conveniently for the space $H_{X}^{\alpha}$ as follows.
@@
Let $0<q_{1}\leq q_{2}<2$ and $\varepsilon$ be any fixed positive number. Let the number $\beta$ be defined as in (2.25). Let $(\xi_{n})_{n\in{\mathbb{N}}}$ be a sequence of increasing positive numbers $\xi_{n}$ satisfying the condition (2.32). Assume that there is a constant $K$ such that for every $v\in H_{X}^{\alpha,1+\varepsilon}$ and every $m\in{\mathbb{N}}_{0}$, $\left\|{P_{m}v}\right\|_{H_{X}^{\alpha,1+\varepsilon}}\leq{\color[rgb]{0,0,0}{K\left\|{v}\right\|_{H_{X}^{\alpha,1+\varepsilon}}}}$. Then for all $n\geq 2$ there exist ${\boldsymbol{y}}_{k,1},...,{\boldsymbol{y}}_{k,n_{k}}\in{\color[rgb]{0,0,0}{U_{0}}}$ and $h_{k,1},...,h_{k,n_{k}}\in L_{2}(U,{\mathbb{C}};\mu)$, $k=0,...,k_{\lceil n/\log n\rceil}$, such that for the operator $\bar{{\mathcal{S}}}^{X}_{n}:={\mathcal{S}}_{\lceil n/\log n\rceil}^{X}$ defined as in (2.29)–(2.31),